OpenStudy (anonymous):

x^2/7 -x = 1/7 use quadratic formula? help me understand this, please? :) Thanks

OpenStudy (nottim):

You know the quadratic formula right?

OpenStudy (anonymous):

yup :) its just the near to last part that screws me up :(

OpenStudy (anonymous):

and the dumb fractions which i detest :/

OpenStudy (nottim):

If you could, If you would, Please show an ywork you've done? My little son?

OpenStudy (nottim):

(Nothing rhymed with star; I went with what I got)

OpenStudy (anonymous):

o.O

OpenStudy (anonymous):

err i dont know what to do with the x^2 :(

OpenStudy (nottim):

Wait...That's x to the exponent 2?

OpenStudy (anonymous):

yes ^^

OpenStudy (anonymous):

You can rewrite it like this \[\frac{1}{7}x^2 - x - \frac{1}{7}\] and then use quadratic formula

OpenStudy (nottim):

I think it just becomes a=1 in the quadratic formula.

OpenStudy (anonymous):

Then you can factor out a 1/7 leaving you with this\[(x^2 - 7x - 1)\]

OpenStudy (nottim):

Can't disagree with that beautiful trinomial.

OpenStudy (nottim):

Can I? I dunno.

OpenStudy (anonymous):

\[x^{2}/7-x=1/7\]

OpenStudy (anonymous):

^^ yes but i moved the 1/7 over to the other side and factored out a 1/7 leaving you with what i wrote above.

OpenStudy (anonymous):

Use the Quadratic Formula to find the solutions of \[x^2 - 7x - 1 = 0\]

OpenStudy (anonymous):

You should get \[\frac{7 \pm 3 \sqrt{5}}{2}\]

OpenStudy (anonymous):

So the final answer will look like the above?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok thanks :)